# Breaking Data Silos: Towards Open and Scalable Mobility Foundation Models via Generative Continual Learning

### v1

@@ Abstract.
Foundation models have revolutionized fields such as natural language processing and computer vision by enabling general-purpose learning across diverse tasks and datasets. However, building analogous models for human mobility remains challenging due to the privacy-sensitive nature of mobility data and the resulting data silos across institutions. To bridge this gap, we propose MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. Without sharing raw data, MoveGCL enables decentralized and progressive model evolution by replaying synthetic trajectories generated from a frozen teacher model, and reinforces knowledge retention through a tailored distillation strategy that mitigates catastrophic forgetting. To address the heterogeneity of mobility patterns, MoveGCL incorporates a Mixture-of-Experts Transformer with a mobility-aware expert routing mechanism, and employs a layer-wise progressive adaptation strategy to stabilize continual updates. Experiments on six real-world urban datasets demonstrate that MoveGCL achieves performance comparable to joint training and significantly outperforms federated learning baselines, while offering strong privacy protection. MoveGCL marks a crucial step toward unlocking foundation models for mobility, offering a practical blueprint for open, scalable, and privacy-preserving model development in the era of foundation models.
Human mobility, foundation models, continual learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/0611footnotetext: Both authors contributed equally to this work.
@@ 1. Introduction
Figure 1. Method comparison with multiple mobility datasets.
@@
  • Extensive experiments demonstrate that MoveGCL achieves performance comparable to joint training with full data access, and significantly outperforms federated learning baselines without data sharing.
@@ 2.2. Mobility Foundation Models
Inspired by the success of foundation models in NLP and CV, recent efforts have explored pre-trained models for urban and mobility domains (Zhou et al., 2024; Zhang et al., 2024; Choudhury et al., 2024). Early studies primarily focused on aggregated mobility data, leveraging mobility flows across cities to build unified spatio-temporal representations, and have demonstrated strong zero-shot transfer capabilities (Li et al., 2024b; Yuan et al., 2024a; Li et al., 2024a). In contrast, individual-level mobility foundation models are less developed. Researchers have explored multi-scale mobility modeling (Yuan et al., 2025a; Zhang et al., 2025; Long et al., 2024), aiming to capture both micro-level behaviors and macro-level patterns essential for generalization. Attempts such as UniTraj (Zhu et al., 2024), TrajBert (Si et al., 2023), and TrajFM (Lin et al., 2024) have explored learning from open trajectory datasets, but these datasets often consist of short-term or non-representative mobility traces that do not reflect regular human movement patterns. As a result, current models struggle to capture the full complexity and diversity of real-world individual mobility. Recently, LLMs have also widely utilized in generating human mobility (Gong et al., 2024; Ju et al., 2025; Feng et al., 2024; JIAWEI et al., 2024), but the gap between natural language and trajectory data suggests that mobility still requires native foundation models, which can later be aligned with LLMs to bridge symbolic reasoning and physical behavior modeling.
@@ 2.3. Continual Learning
Continual learning (Kim et al., 2022; Wang et al., 2024), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge. A central challenge in continual learning is catastrophic forgetting (Li et al., 2019; Wickramasinghe et al., 2023), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks, as seen in approaches like Elastic Weight Consolidation (EWC) and Synaptic Intelligence (SI). Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods, on the other hand, allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
@@ 3.3. Training Pipeline for Mobility Foundation Model Development
where $\theta$ denotes the model parameters, and $\theta_{\mathrm{old},n}$ and $\theta_{\mathrm{new},n}$ correspond to the parameters of $f_{\mathrm{old},n}$ and $f_{\mathrm{new},n}$, respectively. The loss function$\mathcal{L}(\theta;d_{n};f_{\mathrm{old,}n})$ incorporates constraint terms derived from the previous model $f_{\mathrm{old},n}$ to mitigate catastrophic forgetting. Rather than reinitializing from scratch, we optimize $\theta_{\mathrm{new},n}$ starting from $\theta_{\mathrm{old},n}$. Specifically, when $n=0$, we set $f_{\mathrm{old},n}=f_{\mathrm{base}}$; for $n>0$, we have $f_{\mathrm{old},n}=f_{\mathrm{new},\,n-1}$.
@@ 4.2. Model Architecture
The overall location representation $\mathbf{z}_{l}$ is obtained via concatenation ($\oplus$) of these features, and is further transformed by a shared multi-layer perceptron (MLP). This design is sufficient and generalizable because it captures three complementary views of spatial semantics: (1) semantic functions via POI types, (2) actual mobility signals via visitation popularity, and (3) capture where the location sits in the urban layout. Together, they provide a compact yet expressive embedding that generalizes well across different cities with varied spatial structures.
@@
For each input trajectory $x$, we extract a set of mobility behavior descriptive features and encode them into a obility feature descriptor vector $\mathbf{z}_{m}\in\mathbb{R}^{d}$. This feature set comprises: the jump distance $d_{\mathrm{jump}}$ (distance between the current point and the previous point in the trajectory); the waiting time $t_{\mathrm{wait}}$ (time difference between arrivals at the current and previous points in the trajectory); the quantized radius of gyration $r_{\mathrm{gyr}}$; the quantized location entropy $H_{\mathrm{loc}}$; and the city identifier $\mathrm{ID}_{\mathrm{city}}$. These features are embedded via their respective encoders, where $d_{\mathrm{jump}}$ and $t_{\mathrm{wait}}$ are processed by a Transformer-based continuous feature encoder, and $r_{\mathrm{gyr}}$, $H_{\mathrm{loc}}$, and $\mathrm{ID}_{\mathrm{city}}$ are handled by discrete embedding modules. Finally, the five feature embeddings are concatenated to form the mobility behavior vector:
@@
where $\mathbf{z}_{m}$ is the mobility feature descriptor vector defined above; $X_{i}(x)$ denotes the output of the self-attention submodule; $W_{i,r}$ is a learnable projection matrice; $b_{i}$ is the bias term; and $\mathsf{TopK}(\cdot)$ retains only the top $K$ values (corresponding to the $K$ highest-scoring experts), setting the remaining expert scores to $\infty$- ∞ so that their weights after the softmax operation are effectively zero.
@@
(17) $\mathrm{Score}_{\mathrm{similarity}}=\sum_{i=1}^{d}I\cdot E_{\mathrm{DCN},i},$
@@ 5.1. Experimental Settings
We utilize human mobility datasets from multiple cities to evaluate the performance of MoveGCL. Specifically, the datasets cover over eight hundred thousand users and feature a relatively high sampling rate compared to currently available public datasets. Detailed statistics and descriptions are provided in Table 1. For each city, we randomly sample 120,000 trajectories for training, 40,000 for validation, and 40,000 for testing.
Table 1. Basic statistics of mobility data.
City User Trajectory Location
Atlanta 114941 2348218 1175
Chicago 148000 8051522 4166
Los Angeles 161544 16844127 6198
New York 170321 15766369 4988
Seattle 88569 3362353 1046
Washington D.C 134442 11024181 1361
@@
MoveGCL (FullTune) 0.188 0.304 0.125 0.206 0.064 0.114 0.208 0.329 0.199 0.318 0.147 0.257
MoveGCL (ExpertTune) 0.192 0.310 0.132 0.215 0.062 0.108 0.207 0.327 0.195 0.322 0.147 0.259
@@ 5.2. Overall Performance
5.3. Evaluating Order Invariance in Continual Learning
@@ 5.4. Privacy Evaluation
As MoveGCL is built on a generative continual learning framework that does not retain raw data from previous cities, a key question is whether the synthetic data used for replay may inadvertently leak private information from the original training data. To rigorously evaluate the privacy-preserving properties of our approach, following the methodology in (Yuan et al., 2025a, 2024b), we conduct a comprehensive analysis from three complementary perspectives:
@@
  • Membership Inference Attack (Lin et al., 2020a; Shokri et al., 2017): Given a trained model and a set of samples, a membership inference attack assesses whether an classifier can accurately determine which samples were included in the model’s training set based on the model’s outputs.
@@
Following the experimental setup in (Lin et al., 2020b; Shokri et al., 2017), we use the similarity between generated trajectories and their corresponding real input trajectories as the classification feature. For each trajectory $x$, MoveGCL generates $\tilde{x}$ autoregressively conditioned on $x$, and we compute a similarity score $s(x,\tilde{x})$ to form the input to the classifier. The classifier is then tasked with determining whether $x$ was included in the model’s training set. Positive samples consist of real‐world trajectories that were used during training, while negative samples are real trajectories from the same city that were held out. We evaluate the attack success rate, defined as the proportion of samples for which the classifier correctly infers membership status. We employ three widely used classification algorithms: Logistic Regression (LR), Support Vector Machine (SVM) and Random Forest (RF).
@@
[…]
@@ 5.6. Impact of Replay Volume
As shown in Figure 7, the performance on the base cities (WSC) consistently improves with more generated data, demonstrating that replay volume directly influences the model’s ability to retain past knowledge. In contrast, performance on the newly introduced cities (A, L, N) remains largely stable regardless of replay volume, with no consistent trend of improvement or degradation. These results suggest that while synthetic replay is crucial for mitigating forgetting, it has limited effect on new knowledge acquisition. Thus, allocating a moderate amount of generated data offers a practical balance—sufficient to preserve prior knowledge without incurring unnecessary overhead—supporting the long-term scalability of continual mobility modeling.
5.7. In-Depth Analysis
To better understand why MoveGCL is capable of unifying diverse mobility datasets and effectively handling substantial inter-city heterogeneity, we conduct an in-depth analysis of its location embedding layer. Our goal is to examine whether MoveGCL can learn shared spatial representations across cities. To this end, we extract the location embeddings for each city at two stages: (1) after the initial encoder, and (2) after the Deep & Cross Network (DCN). By comparing these two sets of embeddings, we aim to evaluate the role of DCN in aligning spatial semantics across heterogeneous urban environments.
@@ 6. Conclusion
In this work, we present MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. By enabling decentralized model evolution without sharing raw data, MoveGCL addresses key challenges in real-world human mobility modeling, including data silos, privacy constraints, and heterogeneous mobility distributions. MoveGCL represents a significant step toward realizing foundation models for mobility by offering a practical and generalizable framework that facilitates collaborative learning across cities and institutions. It paves the way for long-term, privacy-safe, and adaptive modeling of human movement, with broad implications for urban planning, transportation optimization, and evidence-based policy making.
Despite its promising performance, MoveGCL remains constrained by the limited availability of high-quality training data. The development of more large-scale, semantically rich, and geographically diverse mobility datasets will be crucial for further improving the model’s generalization and robustness. We encourage the broader research community and data-holding institutions to join this collaborative effort, contributing to the creation of open, inclusive, and powerful spatiotemporal foundation models for the mobility domain.

### v2

@@ Abstract.
Foundation models have revolutionized fields such as natural language processing and computer vision by enabling general-purpose learning across diverse tasks and datasets. However, building analogous models for human mobility remains challenging due to the privacy-sensitive nature of mobility data and the resulting data silos across institutions. To bridge this gap, we propose MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. Without sharing raw data, MoveGCL enables decentralized and progressive model evolution by replaying synthetic trajectories generated from a frozen teacher model, and reinforces knowledge retention through a tailored distillation strategy that mitigates catastrophic forgetting. To address the heterogeneity of mobility patterns, MoveGCL incorporates a Mixture-of-Experts Transformer with a mobility-aware expert routing mechanism, and employs a layer-wise progressive adaptation strategy to stabilize continual updates. Experiments on six real-world urban datasets demonstrate that MoveGCL achieves performance comparable to joint training and significantly outperforms federated learning baselines, while offering strong privacy protection. MoveGCL marks a crucial step toward unlocking foundation models for mobility, offering a practical blueprint for open, scalable, and privacy-preserving model development in the era of foundation models. To facilitate reproducibility and future research, we have released the code and models at https://github.com/ScottLiu2003/MoveGCL.
journalyear: 2025copyright: ccconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances inGeographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762728isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Machine learningccs: Information systems Mobile information processing systems11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1. Introduction
Figure 1. Method comparison with multiple mobility datasets.
@@
  • MoveGCL achieves performance comparable to joint training with full data access, and significantly outperforms federated learning baselines without data sharing.
@@ 2.2. Mobility Foundation Models
Inspired by the success of foundation models in NLP and CV, recent efforts have explored pre-trained models for urban and mobility domains (Yuan et al., 2024d; Chai et al., 2025; Han et al., 2025; Zhou et al., 2024; Zhang et al., 2024; Choudhury et al., 2024). Early studies primarily focused on aggregated mobility data, leveraging mobility flows across cities to build unified spatio-temporal representations, and have demonstrated strong zero-shot transfer capabilities (Li et al., 2024b; Yuan et al., 2024b, a; Li et al., 2024a). In contrast, individual-level mobility foundation models are less developed. Researchers have explored multi-scale mobility modeling (Yuan et al., 2025a; Zhang et al., 2025; Long et al., 2024), aiming to capture both micro-level behaviors and macro-level patterns essential for generalization. Attempts such as UniTraj (Zhu et al., 2024), TrajBert (Si et al., 2023), and TrajFM (Lin et al., 2024) have explored learning from open trajectory datasets, but these datasets often consist of short-term or non-representative mobility traces that do not reflect regular human movement patterns. As a result, current models struggle to capture the full complexity and diversity of real-world individual mobility. Recently, LLMs have also widely utilized in generating human mobility (Shao et al., 2024; Gong et al., 2024; Ju et al., 2025; Feng et al., 2024; JIAWEI et al., 2024), but the gap between natural language and trajectory data suggests that mobility still requires native foundation models, which can later be aligned with LLMs to bridge symbolic reasoning and physical behavior modeling.
@@ 2.3. Continual Learning
Continual learning (Kim et al., 2022; Wang et al., 2024, 2024), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge (Yoon et al., 2021). A central challenge in continual learning is catastrophic forgetting (Li et al., 2019; Wickramasinghe et al., 2023), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks, as seen in approaches like Elastic Weight Consolidation (EWC) and Synaptic Intelligence (SI). Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods, on the other hand, allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
@@ 3.3. Training Pipeline for Mobility Foundation Model Development
[…]
@@ 4.2. Model Architecture
The overall location representation $\mathbf{z}_{l}$ is obtained via concatenation ($\oplus$) of these features, and is further transformed by a shared multi-layer perceptron (MLP). This design is sufficient and generalizable because it captures three complementary views of spatial semantics: (1) semantic functions via POI types, (2) actual mobility signals via visitation popularity, and (3) captures where the location sits in the urban layout. Together, they provide a compact yet expressive embedding that generalizes well across different cities with varied spatial structures.
@@
For each input trajectory $x$, we extract a set of mobility behavior descriptive features and encode them into a mobility feature descriptor vector $\mathbf{z}_{m}\in\mathbb{R}^{d}$. This feature set comprises: the jump distance $d_{\mathrm{jump}}$ (distance between the current point and the previous point in the trajectory); the waiting time $t_{\mathrm{wait}}$ (time difference between arrivals at the current and previous points in the trajectory); the quantized radius of gyration $r_{\mathrm{gyr}}$; the quantized location entropy $H_{\mathrm{loc}}$; and the city identifier $\mathrm{ID}_{\mathrm{city}}$. These features are embedded via their respective encoders, where $d_{\mathrm{jump}}$ and $t_{\mathrm{wait}}$ are processed by a Transformer-based continuous feature encoder, and $r_{\mathrm{gyr}}$, $H_{\mathrm{loc}}$, and $\mathrm{ID}_{\mathrm{city}}$ are handled by discrete embedding modules. Finally, the five feature embeddings are concatenated to form the mobility behavior vector:
@@
where $\mathbf{z}_{m}$ is the mobility feature descriptor vector defined above; $X_{i}(x)$ denotes the output of the self-attention submodule; $W_{i,r}$ is a learnable projection matrice; $b_{i}$ is the bias term; and $\mathsf{TopK}(\cdot)$ retains only the top $K$ values (corresponding to the $K$ highest-scoring experts), setting the remaining expert scores to -$\infty$ so that their weights after the softmax operation are effectively zero.
@@
(17) $\mathrm{Score}_{\mathrm{similarity}}=\sum_{i=1}^{d}P\cdot E_{\mathrm{DCN},i},$
@@ 5.1. Experimental Settings
We utilize human mobility datasets from multiple cities to evaluate the performance of MoveGCL. Specifically, the datasets cover over eight hundred thousand users and feature a relatively high sampling rate compared to currently available public datasets. Detailed statistics and descriptions are provided in Appendix Table 4. For each city, we randomly sample 120,000 trajectories for training, 40,000 for validation, and 40,000 for testing.
@@
MoveGCL
0.188 0.304 0.125 0.206 0.064 0.114 0.208 0.329 0.199 0.318 0.147 0.257
MoveGCL
0.192 0.310 0.132 0.215 0.062 0.108 0.207 0.327 0.195 0.322 0.147 0.259
MoveGCL
@@ 5.2. Overall Performance
5.3. Order Invariance in Continual Learning
@@ 5.4. Privacy Evaluation
As MoveGCL is built on a generative continual learning framework that does not retain raw data from previous cities, a key question is whether the synthetic data used for replay may inadvertently leak private information from the original training data. To rigorously evaluate the privacy-preserving properties of our approach, following the methodology in (Yuan et al., 2025a, 2024c), we conduct a comprehensive analysis from three complementary perspectives:
@@
  • Membership Inference Attack (Lin et al., 2020a; Shokri et al., 2017): Given a trained model and a set of samples, it assesses whether an classifier can accurately determine which samples were included in the model’s training set based on the model’s outputs.
@@
Following the experimental setup in (Lin et al., 2020b; Shokri et al., 2017), we use the similarity between generated trajectories and their corresponding real input trajectories as the classification feature. For each trajectory $x$, MoveGCL generates $\tilde{x}$ autoregressively conditioned on $x$, and we compute a similarity score s(x,x~)s($\tilde{x}$) to form the input to the classifier. The classifier is then tasked with determining whether $x$ was included in the model’s training set. Positive samples consist of real‐world trajectories that were used during training, while negative samples are real trajectories from the same city that were held out. We evaluate the attack success rate, defined as the proportion of samples for which the classifier correctly infers membership status. We employ three widely used classification algorithms: Logistic Regression (LR), Support Vector Machine (SVM) and Random Forest (RF).
@@
In our experiment, we randomly select a subset of trajectories and consider two training scenarios: one in which this subset is included in the training data ($D$), and one in which it is excluded ($D^{\prime}$). For each scenario, we train $M$ on the corresponding dataset and then use each selected trajectory as a conditioning input to generate multiple synthetic trajectories. We compute a similarity score between each generated trajectory and its original conditioning trajectory. These similarity scores are modeled as two Gaussian distributions corresponding to $\mathbb{P}\bigl{[}M(z;\,D)\bigr{]}$ and $\mathbb{P}\bigl{[}M(z;\,D^{\prime})\bigr{]}$, respectively. Finally, we estimate the privacy‐budget parameters $\varepsilon$ from these distributions. As shown in Table 3, without applying any additional privacy-preserving mechanisms, MoveGCL naturally achieves a privacy budget of $\varepsilon\approx 1\text{--}2$ for 75% of randomly sampled trajectories. This level is generally considered an acceptable operating point for generative models (Lin et al., 2020a); for example, Apple adopts a privacy budget of $\varepsilon=4.0$ ***https://www.apple.com/privacy/docs/Differential_Privacy_Overview.pdf.
@@ 5.5. Ablation Studies
5.6. In-Depth Analysis
To better understand why MoveGCL is capable of unifying diverse mobility datasets and effectively handling substantial inter-city heterogeneity, we conduct an in-depth analysis of the location embedding layer to examine whether MoveGCL can learn shared spatial representations across cities. To this end, we extract the location embeddings for each city at two stages: (1) after the initial encoder, and (2) after the Deep & Cross Network (DCN). By comparing these two sets of embeddings, we evaluate the role of DCN in aligning spatial semantics across heterogeneous urban environments.
@@ 5.7. Impact of Replay Volume
As shown in Figure 8, the performance on the base cities (WSC) consistently improves with more generated data, indicating that replay volume directly affects the ability to retain past knowledge. In contrast, performance on the newly introduced cities (A, L, N) remains largely stable regardless of replay volume, with no consistent trend of improvement or degradation. These results suggest that while synthetic replay is crucial for mitigating forgetting, it has limited effect on new knowledge acquisition. Thus, allocating a moderate amount of generated data offers a practical balance—sufficient to preserve prior knowledge without incurring unnecessary overhead—supporting the long-term scalability.
@@ 6. Conclusion
In this work, we present MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. By enabling decentralized model evolution without sharing raw data, it addresses key challenges in real-world human mobility modeling, including data silos, privacy constraints, and heterogeneous mobility distributions. MoveGCL represents a significant step toward realizing mobility foundation models by offering a practical and generalizable framework that facilitates collaborative learning across cities and institutions. It paves the way for long-term, privacy-safe, and adaptive modeling of human movement, with broad implications for urban planning, transportation optimization, and evidence-based policy making. The development of more large-scale, semantically rich, and geographically diverse mobility datasets will be crucial for further improving the model’s generalization and robustness. We encourage the broader research community and data-holding institutions to join this collaborative effort, contributing to the creation of open, inclusive, and powerful spatiotemporal foundation models for the mobility domain.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant 2024YFC3307603 and the National Natural Science Foundation of China under 62476152.
Appendix A Dataset
Table 4. Basic statistics of mobility data.
City User Trajectory Location
Atlanta 114941 2348218 1175
Chicago 148000 8051522 4166
Los Angeles 161544 16844127 6198
New York 170321 15766369 4988
Seattle 88569 3362353 1046
Washington D.C 134442 11024181 1361